# UniPLV: Towards Label-Efficient Open-World 3D Scene Understanding by Regional Visual Language Supervision

### v1

@@ 3.3.2 Vision-Point Knowledge Distillation
3.3.3 Vision-Point Matching Learning
We introduce the vision-point matching (VPM) module to further learn fine-grained alignment between images and point clouds. It is a binary classification task that requires the model to predict whether the pixel-point pairs from the projection are positive (matched) or negative (unmatched). VPM mainly includes an attention encoder module and a binary classifier. Given the paring image features and the point cloud features, the image feature acts as the query vector $Q_{I}$, while the point cloud features serve as the key $K_{P}$ and value vectors $V_{P}$. Self-attention is applied to the image features obtaining image attention feature $Q_{attenI}$ Followed cross-attention between the $Q_{attenI}$ and point cloud features as:
@@ 3.3.3 Vision-Point Matching Learning
$ENCODE_{vpm}=\mathcal{FFN}(\mathrm{cat}(\mathrm{\xi}_{1},...,\mathrm{\xi}_{h})),$ (7)
where $\mathcal{FFN}$ is a feed-forward network, which consists of MLPs and activations; After the VPM encoder gets the attention map $G_{attenIP}$ of the projection paring points and pixels, the matching probability is obtained by a binary classifier:
$\hat{f}_{map}=MLP\left(G_{attenIP}\right),\quad f\in R^{r\times 2},$ (8)
where $r$ is the number of points and pixels that matched after projection; $MLP$ denotes the multilayer perceptron for binary classification. The VPM module uses base and novel classes of paring pixels and points for supervision during training.
3.4 Optimization Objective
@@
$\mathbf{M}_{align}$ =$=\begin{cases}1&\text{if }L_{P}==L_{I}\rightarrow Proj(\Gamma_{lc},\mathbf{P}),\\ 0&\text{otherwise}.\end{cases}$
@@
3.5 Multi-Modal and Multi-Task Training
We present a two-stage multi-task training strategy designed to train the multi-modal framework UniPLV.
@@ 3.5 Multi-Modal and Multi-Task Training
Stage 2: Unified multimodal training. The second stage involves joint training of both the image and point cloud branches, employing distinct loss weights to balance their contributions effectively. Throughout the training process, we consistently utilize the AdamW optimizer, chosen for its adaptive learning capabilities and improved convergence properties. The optimizer parameters, specifically the learning rate and weight decay, are set differently for the image and point cloud branch depending on the backbone of each branch. This strategic differentiation in optimizer settings ensures that both branches are training according to their specific network structures and data characteristics, ultimately leading to superior performance in multi-modal training tasks.
Benefits of the two-stage training schemes, particularly for new classes, are detailed in the experimental section.
3.6 Inference
The inference process is depicted in the Figure 2. During inference, we can encode any arbitrary open-vocabulary categories as text queries and compute their similarity with the observed 3D points. Specifically, we associate each point with the category having highest computed cosine similarity. Our method does not require processing images during inference, since we have distilled the image-text alignment to the point cloud.
Methods B12/N3 B10/N5
@@ 3.6 Inference
Table 1: Open-world 3D semantic segmentation comparisons on nuScenes Benchmark. Full supervision: Both base and new classes are gt labels. Open-world: Best open-world results are presented in bold.
Datasets Methods mIoUB mIoUN
Waymo [27] Full Supervision(PTv3) 74.3 81.8
UniPLV(PTv3) 74.0 75.4
SemanticKITTI [1] Full Supervision (PTv3) 57.7 77.4
UniPLV(PTv3) 57.4 62.3
Table 2: Open-world evaluation on the SemanticKITTI and Waymo datasets.
Methods nuScenes
CLIP2Scene [4] 20.8
OpenScenes [22] 42.1
MSeg[11] Voting 31.0
@@
Table 3: Comparison for annotation-free task on nuScenes.
@@
Transfer Pipeline mIoUN
2D-3D projection 28.6
UniPLV 66.5
Table 4: Comparison with projected-only approach.
One-Stage Two-Stage $\mathcal{L}_{logits}$ $\mathcal{L}_{distill_{N}}$ $\mathcal{L}_{distill_{F}}$ $\mathcal{L}_{VPM}$ mIoUB mIoUN
@@
Table 5: Optimization analysis: including ablation study for two stage optimization strategies and various loss functions.
Figure 4: Qualitative results of the annotation-free task.
4 Experiment
4.1 Basic Setups
Datasets. We evaluate our approach on three competitive outdoor datasets: nuScenes [2], SemanticKITTI [1], and Waymo [27]. Following RegionPLC [34], we evaluate the open-world understanding capabilities on two tasks: base-annotated open world (i.e. part of categories annotated) and annotation-free open world (i.e. no category annotated). For the base-annotated open-world task, we split the categories of the mentioned three datasets into base and novel classes, respectively. As for nuScenes, we follow the category partition of RegionPLC, disregarding the ambiguous category ”other_flat” and randomly divide the remaining 15 classes into B12/N3 (i.e. 12 base and 3 novel categories) and B10/N5. We randomly split the category into B15/N2 for Waymo and B17/N2 for SemanticKITTI. Please refer Suppl. for more details.
Implementation Details and Evaluation Metrics. We build the text and image branch upon GLEE [29], adopting CLIP [24] text encoder to process text descriptions and MaskDINO [13] as image encoder-decoder. We utilize the MinkUNet [6], SparseUnet32 [9], and PTv3 [31] as point cloud backbone. We use the ADAM optimizer with a batch size of 32 on 16 NVIDIA A800 GPUs. We use the mean intersection-over-union (mIoU) and harmonic mean IoU (hIoU) as the evaluation indicators. Further training details can be found in the supplementary materials.
4.2 Quantitative Results
4.2.1 Base-Annotated Open-World
Table 1 shows major comparison methods, among which RegionPLC is the current SOTA method. Since the results of public methods are mainly on nuScenes benchmark, we mainly compare with other methods on the nuScenes dataset. In SemanticKITTI and Waymo datasets, only the gap with full supervision and visual analysis are presented.
nuScenes results. For a fair and detailed comparison, we present the results of three point cloud segmentation backbones in Table 1. Using the same point cloud backbone SparseUnet32 as RegionPLC, our method largely lifts the mIoU of unseen categories by 6.2% $\sim$21.8among various partitions on nuScenes without constructing any point-text pairs. When utilizing the higher-performance backbone network, PTv3, our approach achieves a mIou gain of 10.5% $\sim$27.7for the new classes compared to RegionPLC. Even with a lower performance backbone, MinkUNet, our method still outperforms RegionPLC by 0.3% $\sim$20.7for new classes. While achieving significant improvements in the new classes, we can keep the mIOU of the base classes fluctuating within 0.3% $\sim$3compared to fully supervised performance. Comparing the 3DGenZ and 3DTZSL, which omit language alignment, our method even obtains 60.1% $\sim$65.9gain on mIoU of new classes. The results for different backbone substitutability demonstrate the adaptability and integrability of our method.
SemanticKITTI and Waymo results. As shown in Tab 2, partition B15/N2 on Waymo shows only a 6.4% difference in new classes performance compared to the full supervision, which demonstrates the robustness and practicality of our approach in open-world 3D scene understanding. On the SemanticKITTI dataset, the performance of B17/N2 partition differs from fully supervised classes by 15.1%. It should be pointed out that the SemanticKITTI dataset only has forward-looking image, and we only capture point clouds at a forward angle of 120° for experimentation, which leads to a large performance gap with full supervision. The results on these two datasets demonstrate the broad applicability of our proposed method in different scenarios.
4.2.2 Annotation-Free Open-World Segmentation.
@@ 4.2.1 Base-Annotated Open-World
We evaluate the annotation-free capability of our method on the nuScenes validation set, which consists of 16 categories. We freeze the image branch of the proposed UniPLV and load the pre-trained weights from GLEE. Without any 2D or 3D annotations, we have successfully transferred 2D open-world capabilities to 3D through distillation and matching learning. As shown in Table 3, we compare our approach to the most closely related work on zero-shot 3D semantic segmentation: CLIP2Scene [4], OpenScene [22], and MSeg [11] Voting. UniPLV achieves better performance than the above three methods. UniPLV outperforms CLIP2Scene with 36.1%, OpenScene with 14.8%, and MSeg Voting with 25.9%, respectively.
4.3 Qualitative analysis
@@
We further demonstrate the open-world 3D understanding capabilities of the proposed method through visualization. As shown in Figure 3, our method successfully identifies and locates new categories while ensuring the recognition accuracy of base categories. The object detected by the image can be located by the point cloud network, and some false detections in the image can also be adaptively filtered out by the point cloud network. For some relatively small and distant targets with few points, such as traffic cones and distant pedestrians, our method still locates and identifies them. Figure 4 demonstrates the open-world 3D scene understanding ability of our method on annotation-free task. The overall point cloud clustering groups identified by our method are basically consistent with the GT. Background categories, such as plants, drivable areas, sidewalks, and long-tail categories, such as construction vehicles and buses, have been successfully identified. Classes that are easily confused conceptually, such as terrain and sidewalk, may present cross-misrecognition due to the fact that large text models learn a wide range of vocabulary, resulting in blurred boundaries between similar vocabularies.
4.4 Ablation Studies
@@
The ablation experiments are all conducted on nuScenes based on the category partition B12/N3. Multi-Modal Learning Framework. To verify the contrastive learning ability of the proposed framework, we directly project point clouds into images to get 2D pixel semantics predicted by 2D foundation models. As shown in Table 4, our proposed framework outperforms the projection-only approach by 37.9% of mIoU on new classes. This experiment shows that the proposed framework can align the three modes well with limited point clouds data.
@@
Alignment and Optimization. Table 5 presents impact of different modules, corresponding optimization losses, and the two-stage training procedure. In comparative experiments between one-stage and two-stage training, two stage-training yields an improvement of 2.8% $\sim$ 8.3% for both base and new classes. Based on the loss $\mathcal{L}_{logits}$, adding an independent distillation $\mathcal{L}_{distill_{N}}$ for the new class resulted in a 2.5% increase. The feature distillation loss $\mathcal{L}_{distill_{F}}$ from pixels to points results in a 2.6% improvement. The vision-point matching learning loss $\mathcal{L}_{VPM}$ brings 1.3% mIoU gains in the new classes. The experimental results demonstrate the effectiveness of our designed 2D to 3D knowledge transfer alignment strategy and provide a valuable reference for multi-modal fusion tasks involving point clouds and images.
5 Conclusion and Future work
@@
Conclusion. We present a unified multi-modal learning framework, termed UniPLV, for open world 3D scene understanding that does not require point cloud text pairs. Leveraging image as a bridge, we propose logits distillation, feature distillation, and modal matching modules. We further introduce four task-specific losses and a two-stage training procedure for stably multi-modal learning. Our method significantly outperforms SOTA methods on nuScenes. Besides, experimental results on different 3D backbones as well as on Waymo and Semantickitti datasets demonstrate the scalability and lightweight of our method.
Future work. There are some works waiting for improvement and solution in the future. Our proposed framework has been validated exclusively on outdoor datasets. In the future, we aim to extend this validation to indoor datasets, e.g. ScanNet, where the projection parameters between 2D and 3D are more accurate. We will improve and quantify the image branch in the future, so that the proposed framework can achieve both 2D and 3D open-world scene understanding tasks simultaneously. The point cloud branch can also be replaced by an occupancy prediction network to extend the open world application.

### v2

@@ III-C2 Vision-Point Knowledge Distillation
III-C3 Vision-Point Matching Learning
We introduce the vision-point matching (VPM) module to further learn fine-grained alignment between images and point clouds. It is a binary classification task that requires the model to predict whether the pixel-point pairs from the projection are positive (matched) or negative (unmatched). VPM mainly includes an attention encoder module and a binary classifier. Given the paring image features and the point cloud features, the image feature acts as the query vector $Q_{I}$, while the point cloud features serve as the key $K_{P}$ and value vectors $V_{P}$. Self-attention is applied to the image features, obtaining the image attention feature $Q_{attenI}$. Followed cross-attention between the $Q_{attenI}$ and point cloud features as:
@@ III-C3 Vision-Point Matching Learning
$EN_{vpm}=\mathcal{FFN}(\mathrm{cat}(\mathrm{\xi}_{1},...,\mathrm{\xi}_{h})),$ (7)
where $\mathcal{FFN}$ is a feed-forward network, which consists of MLPs and activations. After the VPM encoder gets the attention map $G_{attenIP}$ of the projection paring points and pixels, the matching probability is obtained by a binary classifier:
$\hat{f}_{map}=\operatorname{MLP}\left(G_{attenIP}\right),\quad f\in R^{r\times 2},$ (8)
where $r$ is the number of points and pixels that matched after projection; $\operatorname{MLP}$ denotes the multilayer perceptron for binary classification. The VPM module uses base and novel classes of pairing pixels and points for supervision during training.
III-D Optimization Objective
@@
[…]
@@
III-E Multimodal Training and Inference
We present a two-stage multi-task training strategy designed to train the multimodal framework UniPLV.
@@ III-E Multimodal Training and Inference
Stage 2: Unified multimodal training. The second stage involves joint training of both the image and point cloud branches, employing distinct loss weights to balance their contributions effectively. Throughout the training process, we consistently utilize the AdamW optimizer, chosen for its adaptive learning capabilities and improved convergence properties. The optimizer parameters, specifically the learning rate and weight decay, are set differently for the image and point cloud branches depending on the backbone of each branch. This strategic differentiation in optimizer settings ensures that both branches are training according to their specific network structures and data characteristics, ultimately leading to superior performance in multimodal training tasks. The benefits of the two-stage training schemes, particularly for new classes, are detailed in the experimental section.
The inference process is depicted in the Fig. 2. During inference, we can encode any arbitrary open-vocabulary categories as text queries and compute their similarity with the observed 3D points. Specifically, we associate each point with the category having the highest computed cosine similarity. Our method does not require processing images during inference, since we have distilled the image-text alignment to the point cloud.
IV Experiment
TABLE I: Open-world 3D semantic segmentation comparisons on nuScenes Benchmark. Full supervision: Both base and new classes are gt labels. Open-world: Best open-world results are presented in bold.
Type Method B12/N3 B10/N5
@@ IV Experiment
TABLE II: Open-world 3D semantic segmentation comparisons on ScanNet. Full Supervision: Both base and new classes are gt labels. Open-World: Best open world results are presented in bold.
Type Method B15/N4 B12/N7 B10/N9
mIoUB mIoUN mIoUB mIoUN mIoUB mIoUN
Full Supervision SparseUnet32 [24] 68.4 79.6 71.2 71.6 76.5 71.9
Open-World 3DGenZ [27] 56.0 12.6 35.5 13.3 63.6 6.6
3DTZSL [28] 36.7 6.1 36.6 2.0 55.5 4.2
Ovseg-3D [29] 64.4 0.0 55.7 0.1 68.4 0.9
PLA [5] 68.3 62.4 69.5 45.9 76.2 40.8
RegionPLC [22] (SparseUnet32) 68.2 70.7 69.9 66.6 76.3 55.6
UniPLV(SparseUnet32) 68.4 75.7 69.9 69.8 76.6 67.2
IV-A Basic Setups
Datasets. We evaluate our approach on three competitive outdoor datasets: nuScenes [30], SemanticKITTI [31], and Waymo [32] as well as the indoor dataset ScanNet [33]. Following RegionPLC [22], we evaluate the open-world understanding capabilities on two tasks: base-annotated open world (i.e., part of categories annotated) and annotation-free open world (i.e., no category annotated). For the base-annotated open-world task, we split the categories of the mentioned three datasets into base and novel classes, respectively. As for nuScenes, we follow the category partition of RegionPLC, disregarding the ambiguous category ”other_flat” and randomly dividing the remaining 15 classes into B12/N3 (i.e., 12 base and 3 novel categories) and B10/N5. We randomly split the category into B15/N2 for Waymo and B17/N2 for SemanticKITTI. Regarding the ScanNet dataset, we adopt the category partition of PLA [5].
TABLE III: Open-world evaluation on the SemanticKITTI and Waymo datasets.
[…]
Waymo Full Supervision 74.3 81.8
UniPLV 74.0 75.4
SemanticKITTI Full Supervision 57.7 77.4
UniPLV 57.4 62.3
Implementation Details and Evaluation Metrics. We build the text and image branch upon GLEE [7], adopting CLIP [17] text encoder to process text descriptions and MaskDINO [34] as image encoder-decoder. We utilize the MinkUNet [25], SparseUnet32 [24], and PTv3 [26] as point cloud backbone. We use the ADAM optimizer with a batch size of 32 on 16 NVIDIA A800 GPUs. We use the mean intersection-over-union (mIoU) and harmonic mean IoU (hIoU) as the evaluation indicators.
IV-B Quantitative Results
IV-B1 Base-Annotated Open-World
Table I shows major comparison methods, among which RegionPLC is the current SOTA method. Since most public results are reported on nuScenes, the primary comparison focuses on this benchmark. For SemanticKITTI and Waymo, we present the gap with full supervision and visual analysis.
nuScenes Results. For a fair and detailed comparison, we present the results of three point cloud segmentation backbones in Table I. Using the same point cloud backbone, SparseUnet32 as RegionPLC, our method largely lifts the mIoU of unseen categories by 6.2% $\sim$21.8% among various partitions on nuScenes without constructing any point-text pairs. When utilizing the higher-performance backbone network, PTv3, our approach achieves a mIoU gain of 10.5% $\sim$27.7% for the new classes compared to RegionPLC. Even with a lower performance backbone, MinkUNet, our method still outperforms RegionPLC by 0.3% $\sim$20.7% for new classes. While achieving significant improvements in the new classes, we can keep the mIOU of the base classes fluctuating within 0.3% $\sim$3% compared to fully supervised performance. Comparing the 3DGenZ and 3DTZSL, which omit language alignment, our method even obtains 60.1% $\sim$65.9% gain on mIoU of new classes. The results for different backbone substitutability demonstrate the adaptability and integrability of our method.
TABLE IV: Comparisons for annotation-free task on nuScenes.
[…]
nuScenes CLIP2Scene 20.8
OpenScenes 42.1
MSeg Voting 31.0
2D-3D projection 28.6
@@ IV-B1 Base-Annotated Open-World
ScanNet Results. As shown in Table II, UniPLV outperforms the previous state-of-the-art method, RegionPLC. Specifically, in the performance on unseen categories, our approach exceeds RegionPLC by a margin of 3.2% to 11.6%, while the gap relative to fully supervised methods is only 1.8% to 4.7%.
SemanticKITTI and Waymo Results. As shown in Tab III, partition B15/N2 on Waymo shows only a 6.4% difference in new classes performance compared to the full supervision, which demonstrates the robustness and practicality of our approach in open-world 3D scene understanding. On the SemanticKITTI dataset, the performance of the B17/N2 partition differs from fully supervised classes by 15.1%. It should be pointed out that the SemanticKITTI dataset only has forward-looking images, and we only capture point clouds at a forward angle of $120^{\circ}$ for experimentation, which leads to a large performance gap with full supervision. The results on these two datasets demonstrate the broad applicability of our proposed method in different scenarios.
@@
TABLE V: Optimization analysis: including ablation study for two-stage optimization strategies and various loss functions.
[…]
@@
IV-B2 Annotation-Free Open-World Segmentation.
@@
We evaluate the annotation-free capability of our method on the nuScenes validation set, which consists of 16 categories. We freeze the image branch of the proposed UniPLV and load the pre-trained weights from GLEE. Without any 2D or 3D annotations, we have successfully transferred 2D open-world capabilities to 3D through distillation and matching learning. As shown in Table IV, we compare our approach to the most closely related work on zero-shot 3D semantic segmentation: CLIP2Scene [18], OpenScene [2], and MSeg [35] Voting. UniPLV achieves better performance than the above three methods. UniPLV outperforms CLIP2Scene with 36.1%, OpenScene with 14.8%, and MSeg Voting with 25.9%, respectively.
IV-C Qualitative analysis
@@
We further demonstrate the open-world 3D understanding capabilities of the proposed method through visualization. As shown in Fig. 3, our method successfully identifies and locates new categories while ensuring the recognition accuracy of base categories. The object detected by the image can be located by the point cloud network, and some false detections in the image can also be adaptively filtered out by the point cloud network. For some relatively small and distant targets with few points, such as traffic cones and distant pedestrians, our method still locates and identifies them.
IV-D Ablation Studies
@@
The ablation experiments are all conducted on nuScenes based on the category partition B12/N3. Multimodal Learning Framework. To verify the contrastive learning ability of the proposed framework, we directly project point clouds into images to get 2D pixel semantics predicted by 2D foundation models. As shown in Table IV, our proposed framework outperforms the projection-only approach by 37.9% of mIoU on new classes. This experiment shows that the proposed framework can align the three modes well with limited point cloud data.
@@
Alignment and Optimization. Table V presents the impact of different modules, corresponding optimization losses, and the two-stage training procedure. In comparative experiments between one-stage and two-stage training, two-stage training yields an improvement of 2.8% sim\ sim8.3% for both base and new classes. Based on the loss $\mathcal{L}_{logits}$, adding an independent distillation $\mathcal{L}_{distill_{N}}$ for the new class resulted in a 2.5% increase. The feature distillation loss $\mathcal{L}_{distill_{F}}$ from pixels to points results in a 2.6% improvement. The vision-point matching learning loss $\mathcal{L}_{VPM}$ brings 1.3% mIoU gains in the new classes. The experimental results demonstrate the effectiveness of our designed 2D to 3D knowledge transfer alignment strategy and provide a valuable reference for multimodal fusion tasks involving point clouds and images.
V Conclusion and Future Work
@@
We present a unified multimodal learning framework, termed UniPLV, for open-world 3D scene understanding that does not require point cloud text pairs. Leveraging image as a bridge, we propose logits distillation, feature distillation, and modal matching modules. We further introduce four task-specific losses and a two-stage training procedure for stably multimodal learning. Our method significantly outperforms SOTA methods on nuScenes and ScanNet. Besides, experimental results on different 3D backbones as well as on Waymo and SemanticKITTI datasets demonstrate the scalability and lightweight of our method. We will improve and quantify the image branch in the future, so that the proposed framework can achieve both 2D and 3D open-world scene understanding tasks . The point cloud branch can also be replaced by an occupancy prediction network.